Two conductors are made of the same material and have the same length. Conductor A is a

solid wire of diameter 1.0 mm. Conductor B is a hollow tube of outside diameter 2.0 mm

and inside diameter 1.0 mm. What is the resistance ratio RA/RB, measured between their

ends?


1
Expert's answer
2021-06-03T18:26:35-0400

The resistance in terms of the area and the length of the wire is given by:

"R = \frac{ρL}{A}"

if we have two wires, the first one is a solid wire with a diameter of "d_A=1 \times 10^{-3} \;m" , and the second one is a hollow wire with inner diameter of "d_{B,i}=1 \times 10^{-3} \; m" and outer diameter of "d_{B,o}=2 \times 10^{-3} \;m" , so the cross sectional area of the first wire is:

"A_A = \pi R^2_A = \frac{\pi d^2_A}{4}"

Hence the resistance is:

"R_A = \frac{4ρL_A}{\pi d^2_A}"

The area of the second wire is:

"A_B = \pi r^2_{B,o} - \pi r^2_{B,i} = \frac{\pi}{4}(d^2_{B,o} -d^2_{B,i})"

Hence the resistance is:

"R_B = \frac{4ρL_B}{\pi(d^2_{B,o} -d^2_{B,i})}"



The ratio between the resistances:

"\frac{R_A}{R_B} = \frac{(d^2_{B,o} -d^2_{B,i})L_A}{d^2_AL_B}"

The wires have the same length, therefore:

"\frac{R_A}{R_B} = \frac{d^2_{B,o} -d^2_{B,i}}{d^2_A} \\

= \frac{(2 \times 10^{-3})^2 -(1 \times 10^{-3})^2}{(1 \times 10^{-3})^2} = 3 \\

\frac{R_A}{R_B} = 3"